OpenStudy (anonymous):

Use the definition f'(a)=lim h->0 (f(a+h)-f(a))/h to find the derivative of f(x)=(1/x) at x=3. Please be detailed so I can clearly see the steps I can use for similar problems.

OpenStudy (anonymous):

find f(x+h)= 1/(x+h), when x=3 , this works out to 1/(3+h) f(x+h)-f(x)=1/(3+h)-1/3 that should give you a hint

OpenStudy (anonymous):

Well, that is what I did. Then I got ((1/3+h)-(1/3))/h The I simplified that to (1/3 + 1/h -1/3)/h , the 1/3 -1/3 cancel out, leaving (1/h)/h or (1/h) x (1/h) = (1/h)^2. What I am confused about is h -> 0, so now I plug 0 into (1/h)^2, and I know you cannot devide by 0, which makes me think I did something wrong or that the answer is simply 0. Am I correct? If not could you please show me the remaining steps, please?

OpenStudy (anonymous):

You cant write 1/(3+h) = 1/3 + 1/h

OpenStudy (anonymous):

thats the reason yu are getting errors

OpenStudy (anonymous):

But even if I don't write 1/(3+h) as 1/3+1/h, I still have to plug in 0 into the equation, right? So ((1/3+h)-1/3)/h = (1/3-1/3)/0 = 0/0. Would that be 0 or would that be 1? Since anything divided by anything is 1? Or am I still wrong?

OpenStudy (anonymous):

*What I meant by anything / anything is a number divided by itself.

OpenStudy (anonymous):

1/(3+h) - 1/3=(3-3-h)/3(3+h) so you have (-h)/(3)(3+h) now your limit changes to lim -h ------------- h->0 3*(3+h)*h

OpenStudy (anonymous):

check the steps i have suggested, pardon me for any errors

OpenStudy (anonymous):

I'm sorry, I'm confused. Is there any easier way? :(

OpenStudy (anonymous):

unfortunately NO! , the reason is you need to use the limit definition of derivative

OpenStudy (anonymous):

Could you do the whole problem through so I can see how to properly do it? This problem is limiting me from finishing the rest of my homework and correctly understand what I'm supposed to do.

OpenStudy (anonymous):

cancel out the -h and the h in denominator you are left with -1/ (9+3h) , now evaluate the limits put h=0, you get -1/9 as the answer which IS THE ONLY CORRECT ANSWER alternatively if you want to verfiy the derivative of 1/x the answer of that using the x^n formula is -1/x^2 which verifies the work

OpenStudy (anonymous):

let me look this through

OpenStudy (anonymous):

there is no mistake, the only thing i am worried is text appearing as garbled due to unicode format

OpenStudy (anonymous):

text if fine, i just need my brain to understand it now. ill let you know if i have any more questions :)

OpenStudy (anonymous):

*text is fine

OpenStudy (anonymous):

Take your time